Suppose the government implements a price ceiling of $20/unit in this market.  Is this price ceiling binding on the market?

QS = QD

100 - 2(P - 20) = -20 + P

160 = 4P

P = 40

Therefore QD = -20 + 2(40) = 60 units


  1. What is the value of the economic surplus that would be generated in the original equilibrium? Is there a deadweight loss due to the price ceiling policy, and if so, what is its value? Briefly explain. 




1
Expert's answer
2021-10-07T16:50:34-0400

At equilibrium price, consumer surplus = "\frac{1}{2} \times base \times height"

"= \frac{1}{2} \times 40 \times (100 - 30) \\
=20 \times 70 = 1400\\
= 1400"

Economic surplus = Consumer Surplus + Producer Surplus

Deadweight loss generated due to the price ceiling is: "(\frac{1}{2} \times base \times height) \times 2"

"=(\frac{1}{2} \times (40 - 20) \times (30 -20)) \times 2 \\
= 20 \times 10\\
= 200"